x(1+y) 1/2 +y(1+x) 1/2 and x is not equal to y, then dy/dx = ???

x(1+y)1/2+y(1+x)1/2 and x is not equal to y, then dy/dx = ???

Grade:10

1 Answers

Arun
5 years ago
Dear student
 
x * ½ (1 + y)^-1/2 y’ + (1 + y)^1/2 + y * ½ (1 + x)^-1/2 + (1 + x)^1/2 * y’ = 0
 
on simplifying you get
 
y’ = dy/dx = 1 / (x + 1)2